# Quantformer: from attention to profit with a quantitative transformer trading strategy

### v2

@@ 3.6 Prediction
The Mean Squared Error Loss (MSELoss) will be used to quantify the loss. MSE Loss is a widely used metric for assessing the discrepancy between a model’s predictions and the actual values. It is defined as the average of the squares of the differences between the predicted value and the actual prices:
$\mathrm{MSELoss}=\frac{1}{N}\sum^{N}_{i=1}(y_{i}-\hat{y}_{i})^{2}$ (8)
In equation (8), $y_{i}$ is the actual value of the $i^{\mathrm{th}}$ observation, $\hat{y}_{i}$ is the corresponding model prediction, and $N$ is the total number of input samples. During each epoch, the MSE Loss is computed to guide the optimization process, to minimize this loss over successive iterations. A lower MSE indicates a closer alignment between the model’s predictions and the actual stock prices, signifying an improvement in the model’s learning and its ability to generalize from the training data.
@@ 4 Experiments
Based on the methodology in Section 3, in this section, the set of experiments is introduced, including the data resource, implementation details, trading strategy and metrics.
@@ 4.1 Dataset
The training data of the quantformer comes from the Chinese exchange market. The dataset is collected from the Shanghai Stock Exchange (SHSE) and Shenzhen Stock Exchange (SZSE), which contains 4601 stocks that are listed or had been listed from January 2010 to May 2023. The data comes from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which are quantitative finance terminals. The training period is from January 2010 to December 2019 and the testing period starts from January 2020.
@@
Closing price adjustments are used for the stock training price [48]. Closing price adjustments are essential in stock market analysis, particularly when analyzing historical data for long-term trends and patterns. This price takes into account factors such as dividends, stock splits, and other corporate actions that can affect a stock’s price over time [49]. By adjusting for these events, the adjusted closing price provides a more accurate picture of a stock’s value and performance, which is more appropriate for financial backtesting.
4.2 Timestamp
@@
Data frequency concerns the number of data points within a specific unit of time, which may reflect different market characteristics [50]. To better test the performance of the factor, three frequencies are considered: monthly, weekly, and daily stock data.
@@
In the first experiments, a timestamp was set as one month. So for a stock $s_{i}$ from the stock set $S^{t}$, it details a feature sequence in which each item in the sequence contains the accumulated profit and accumulated turnover rate in a month. Both the accumulated profit and accumulated turnover rate are the sum of the trading day’s data in that month. In some cases, if one stock was first listed on the market or resumed trading in one month, the available data may not cover whole trading days, the data of the stock from that from would still be recorded. However, if there is a whole month during which the stock did not have any trading (perhaps due to stopped trading or not being listed in that month), this month’s data will be recorded as “NaN”. In the sequence $\chi^{t}_{k}$, the sequence will not be used to train the model when there are missing values in the sequence.
Within the first experiments, three sub-experiments were set. The sub-experiments share the same inputs. For output data, outputs with different dimensions and different lengths but the same dimension are used. $n=5$ and $n=7$ are used in this experiments. In the group of $n=5$, the first one deleted the data with output such as $[0,0,0]^{T}$, and the second one used the origin dataset (zero-output).
@@
Outliers of both accumulated profit and accumulated turnover rate will not be removed from the dataset these situations may happen in the future and they are expected to be predicted, though these incidents rarely happen.
@@
In the second and third experiments, the inputs are in weekly and daily frequency, respectively. The sub-experiments under different groups are similar to the first one. All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and dimension are shown in Table 1.
Table 1: Detailed information of experiments
@@
Strategy Frequency Output Dim Training samples Section zero-output
@@ 4.3 Implementation details
The training process and model are implemented with PyTorch [51]. The parameters of the model are optimized by grid search, which is reliable in low dimensional spaces and simple to implement [52]. Here the input dimension is 2, and the dimension of the hidden feature space $d_{f}$ is 16. The number of multi-head attention modules is 16, and the number of layers of encoder and decoder is 6. The model was trained on an NVIDIA GeForce RTX 2080 GPU and NVIDIA A100 Tensor Core GPU using the Adam optimizer [53] for 50 epochs, the learning rate is 0.001, and the batch size is 64.
@@ 4.4 Trading strategy
A trading strategy is used for the model and Algorithm 1 shows the pseudocode of the strategy. Before the first trade date of the timestamp $t$, all sequence $\chi^{t}_{i,k}$ from the stock set $S^{t}$ will be put in the model and obtain the list of outputs. Then, the stocks will be ranked according to the first element of the output and the first $\frac{1}{q}\%$ stocks will be added to the stock pool. If the stock already was in the stock pool on the last timestamp, it will be held; if the stock is in the predicted pool but not in the previous pool, it will be bought in with the same proportion of the whole account. Stocks that are not in the predicted pool will be sold out. The same method is run repeatedly during the subsequent periods. The backtest starts from January 2020, in other words, the result of the sequences from May 2018 to December 2019 will be used as the first stock pool to trade.
@@
1:The sequence $x^{t}_{i,k}$ for each stock $s_{i}$ in the set $S^{t}$ and the amount of cash.
2:run_monthly(trade, monthday=1, time=‘open’)
3:function obtain_stockpool($x^{t}_{i,k}$):
4:     stockpool $=$ sort(model($x^{t}_{k}$), key=lambda x : x[1], reverse=True)
5:     return stockpool
@@
7:function trade
8:     stockpool $=$ obtain_stockpool($x^{t}_{i,k}$)
9:     for stock in previous_stockpool do
10:         if stock not in stockpool then
11:              order(stock, 0) $\triangleright$ sellout
12:         end if
13:     end for
14:     for stock in stockpool do
15:         if stock not in previous_stockpool then
16:              order(stock, amount / len(stockpool)) $\triangleright$ buy
17:         end if
18:     end for
19:end function
@@ 4.5 Metrics
Formally, the Sharpe Ratio (SR) and the $\alpha$ ratio will be used to test the performance of the strategy. The SR [3] is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase, which is shown in equation (9), where $R_{p}$ is the return of the portfolio and $R_{f}$ is the risk-free rate is the rate of London Interbank Offered Rate (LIBOR) which calculated averaged from estimates submitted by banks in London. The risk-free rate in the backtest will be calculated under the average LIBOR rates during the test period.
@@
Alpha represents the excess return of an investment portfolio relative to its benchmark. It reflects the stock selection skills of an investment. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment. Specifically, alpha is the excess return of the actual portfolio return over its expected theoretical return. Equation (10) shows the calculation method of the alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and systemic risk.
$Alpha=(R_{p}-R_{f})-\beta(E[R_{m}]-R_{f})$ (10)
Besides the SR ratio and alpha rate, the annual return (AR) of the stock, the annual excess return (AER), the average turnover rate (TR) of the portfolio, and the win rate (WR) will also shown. The excess returns are returns achieved above and beyond the return of a proxy and the CSI 300 index is used as the basis when calculating the excess returns. The turnover rate is the average percentage of the portfolio adjusted daily in terms of market. The win rate is the percentage of trade days where the portfolio encounters a positive return overall trade days.
$VaR=\mu+\sigma N^{-1}(X)$ (11)
@@ 5 Results and discussion
Building upon the analysis of individual strategies, this section presents an evaluation and discussion of the experimental outcomes. The discussion focuses on the comparative effectiveness of different timestamps (monthly, weekly, and daily) and different training scales. The results underscore the nuanced relationship between trading frequency and market performance, offering valuable insights for the development of robust quantitative trading strategies.
@@ 5.1 Overall performance
The result of the experiment is shown in Table 2. For the monthly group, Month_1 shows an annual return of 17.35% and an annual excess return of 19.43%, indicating a robust performance over that period. Conversely, Month_3 shows a diminished annual return and annual excess return of 7.37% and 9.91% respectively. The turnover rate is notably high in the second month at 51.69%. The win rate ranges from 43.1% to 57.3%, with the first month showing the highest win rate. The Sharpe Ratio and Alpha, vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values. In the comparison of 99% VaR, the strategies in monthly frequency perform better than others, where Month_1 and Month_3 are lower than 3%, which means the portfolio maximum loss 2.81% and 2.3% of return a day in 99% confidence.
@@
In the weekly strategy results, Week_3 showed a remarkable annual return of 12.3% and an annual excess return of 12.73%, while Week_1 displayed a negative annual return. For the daily strategies, Day_2 had a return of 10.23% and an annual excess return of 10.94%. The turnover rate across the daily strategies showed lesser variation than the monthly or weekly, indicating a more consistent trading frequency. The win rate for the daily strategies remained fairly stable, hovering around the mid-40% mark. The Sharpe Ratio and Alpha for weekly and daily strategies show a mix of positive and negative values, reflecting the fluctuating nature of shorter-term trading efficacy.
To demonstrate the advantages of the quantformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^3] https://www.joinquant.com/ (one of China’s largest financial quantitative platforms) were selected for backtesting under the same trading strategy (Algorithm 1). Price-volume factors were used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by quantformer factors (also used stock turnover rates and price data). The detailed performance of factors and benchmark is shown in Appendix A. On average, these factors achieved an annual return of -3.78%, an average excess return of -2.15%, and a Sharpe ratio of -0.36, where the benchmark return is -1.77%. The average maximum drawdown of the strategy was 44.88%, which means the potential worst-case scenario or the most extreme possible loss. The average volatility of the strategy ($\sigma_{p}$) was 0.26, with the benchmark at 0.197.
@@
Among these factor-based strategies, the quantformer factor performed well. Except for the win rate, which was slightly lower than a few strategies, it excelled in terms of annual return, annual excess return, $\sigma_{p}$, Sharpe Ratio and sortino ratio ranking the best among the 101 factors. The 99% VaR of quantformer factor is in the quantile of the first 10% among all factors. Figure 5 illustrates the return curves of the quantformer (QF_Month_1) alongside the benchmark and some traditional factors. “25_week_rank” represents the current price’s position over the past 25 weeks; “EMAC20” and “MAC20” are the 20-day index moving average and the stock’s 20-day moving average respectively; “DAVOL20” and “VOL20” represent the ratio of the 20-day average turnover rate to the 120-day average turnover rate, and the mean of the stock’s 20-day turnover rate; “ROC20” is the price rate of change over 20 days; “MAWVAD” is calculated as the product of the difference between the closing price and the opening price, divided by the range of the highest and lowest prices, all multiplied by the volume, accumulated over six days; “Variance20” is the variance of the stock’s 20-day annualized returns. Most of these factors are compared based on a window of 20 timestamps, as the quantformer factor was also trained on 20 timestamps. In comparison with these factors, the quantformer factor (blue line) demonstrated a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies, expressing the improvement of the quantformer factor over traditional price-volume factors.
5.2 Half-life period test
To test the half-life period of the factor, another three portfolios based on $Month\_3$ were conducted. The first one is the stock pool used in the next month; the second is using it in the month after the month; the last one is using it in the third month after the last feature of inputs. For instance, in the first strategy, the trade will start in January 2020; in the second, it will start in February 2020; in the second, it will start in March 2020. For the period from January 2020 to February 2022, the strategy achieved a total return (TotalR) of 65.29% and an annual return of 17.87%. The total excess return (TER) over this period was 66.39%, indicating a strong performance above a benchmark or risk-free rate. The Sharpe Ratio for this period was 0.863, which is a respectable figure that suggests a favorable risk-adjusted return. Additionally, the strategy generated an alpha of 0.164, demonstrating its ability to outperform the expected return on a risk-adjusted basis.
Table 3: Result of the same strategy with different starting month
Strategy TotalR AR TER SR Alpha CSI300
2020/01-2022/02 65.29% 17.87% 66.39% 0.863 0.164 -0.66%
2020/02-2022/03 67.52% 18.21% 65.58% 0.940 0.161 1.17%
2020/03-2022/04 56.41% 15.63% 52.96% 0.739 0.135 2.26%
In the subsequent period from February 2020 to March 2023, the strategy yielded a slightly higher total return of 67.52% and an annual return of 18.21%. The excess return was 65.58%, with an improved Sharpe Ratio of 0.940, reflecting a better risk-adjusted return compared to the previous period. The alpha for this interval was 0.161, indicating the strategy’s consistent performance in delivering additional returns.
Lastly, Table 3 shows the result of the backtest. From March 2020 to April 2023, the strategy’s total return decreased to 56.41%, with an annual return of 15.63% and an excess return of 52.96%. The Sharpe Ratio also declined to 0.739, but it still represents a relatively efficient investment compared to the volatility of returns. The alpha reduced to 0.135, which, while lower than the other periods, still indicates that the strategy managed to provide returns above those predicted by its risk profile.
Comparing the results between monthly, weekly, and daily strategies provides further insights. The monthly strategies can deliver the highest returns, suggesting that longer-term signals better capture short-term trends. The weekly strategies underperform monthly ones but outperform the daily frequency overall. This indicates weekly data strikes a balance between filtering noise and responding swiftly to emerging patterns.
The daily frequency strategies perform lower annual returns and higher risk compared to others This highlights the challenge of making profitable trades at higher frequencies - noise and volatility make reliable signals more difficult. The distribution of returns also shows increased dispersion at the daily level. However, daily strategies achieve more consistent turnover rates around 5-7%. This allows for portfolio adjustments while avoiding excessive trading costs. In contrast, monthly strategies see a turnover rate above 50% in some cases. There is likely an optimal point between 1 week to 1 month where matches market regularly.

### v3

@@ 3.5 Prediction
The Mean Squared Error Loss (MSELoss) is used to quantify the loss. It is defined as the average of the squares of the differences between the predicted value and the actual prices:
[…]
where $||\cdot||^{2}_{2}$ denotes the squared Euclidean distance between predicted and target class probability vectors.
@@ 4 Experiments
In this section, we test the ability of the quantformer architecture, introduced in the previous section, for stock price prediction. We are going to detail the set of experiments, including the data resource, implementation details, trading strategy and metrics.
@@ 4.1 Dataset
Our training data contains 4601 stocks listed either on the Shanghai Stock Exchange (SHSE) or on the Shenzhen Stock Exchange (SZSE). The time period ranges from January 2010 to May 2023. The data has been obtained from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which are quantitative finance terminals. The training period is from January 2010 to December 2019 and the testing period starts from January 2020.
@@
Closing price adjustments, such as dividends, stock splits, and other corporate actions that can affect a stock’s price, are applied to the stock training prices. These adjustments are essential for stock price analysis, especially on long time period, as they provide a more accurate picture of a stock’s value and performance over time, which is more appropriate for financial backtesting. (Diamond, 1971, Wei et al., 2022).
4.2 Time frequency
@@
Data frequency refers to the number of data points within a specific unit of time, which may reflect different market characteristics (De Prado, 2018). We consider three frequencies: monthly, weekly, and daily stock data.
@@
In the first experiments, the time frequency is set to one month. For a stock $s^{t}_{n}$ from the stock set $S^{t}$, each item in the feature sequence contains the accumulated profit and accumulated turnover rate in the month (equation (3)). Sometimes, the available data may not cover whole trading days, for example, because a stock was first listed on the market or resumed trading after an interruption. In such a situation, we would still record the stock data for this day. However, if there is a whole month during which the stock did not have any trading (perhaps due to stopped trading or not being listed in that month), this month’s data will be recorded as “NaN”, and the sequence $\mathcal{X}^{t}_{n}$ will not be used to train the model.
Within the first experiment, three sub-experiments are set. The sub-experiments share the same inputs. The values $\varrho=3$ and $\varrho=5$ are used in this experiment. The value of $\varphi$ is fixed for all the experiments to 0.2. In the groups of $\varrho=5$, stocks are divided into five parts and marked respectively. Then, in the groups of $\varrho=3$, the top $20\%$, middle $20\%$ and bottom $20\%$ of the stocks are labeled with one-hot vectors $[1,0,0],[0,1,0]$ and $[0,0,1]$, respectively. Stocks in the intermediate quantile (i.e, $20\%-40\%$ and $60\%-80\%$) are marked as $[0,0,0]$ (as the boundary term $\xi=0.2$ here, and $\varphi=0.2$ as well). In the first group of $\varrho=3$, the null labels ($0,0,0]^{T}$) are excluded from training, whereas the second group of $\varrho=3$ includes them.
@@
Outliers of both accumulated profit and accumulated turnover rate are not removed from the dataset. These situations may happen in the future and they are expected to be predicted, though these incidents rarely happen.
@@
In the second and third experiments, the inputs are in weekly and daily frequency, respectively. The sub-experiments under different groups are similar to the first one. All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and output dimensions are shown in Table 2.
Table 2: Detailed information about experiments
@@
Strategy Frequency Lable dim($\varrho$) Training samples Section Null-label
@@ 4.3 Implementation details
The model and its training process are implemented with PyTorch (Paszke et al., 2019). The hyperparameters of the model are optimized by grid search, as it is simple to implement and reliable in low-dimensional spaces (Bergstra and Bengio, 2012). Here the input dimension is 2, and the dimension of the hidden feature space $d$ is 16. The number of multi-head attention modules is 16, and the number of layers of encoder and decoder is 6. The model was trained on an NVIDIA GeForce RTX 2070 GPU and NVIDIA A100 Tensor Core GPU using the Adam optimizer (Kingma and Ba, 2015) for 50 epochs. The learning rate is 0.001, and the batch size is 64.
@@ 4.4 Trading strategy
Algorithm 1 shows the pseudocode of the trading strategy. Before the first trade date of the timestamp $t$, all sequences $\mathcal{X}^{t}_{n}$ from the stock set $S^{t}$ are put in the model and the list of outputs $\hat{Y}^{t}$ is required. The portfolio value at time $t$, denoted as $P^{t}$ is updated based on the previous period’s weights and returns as follows:
$P^{t}=P^{t-1}(\sum^{N}_{n=1}w^{t-1}_{n}(1+r^{t-1}_{n}))$ (14)
where $w^{t}_{n}$ is the weight of stock $s^{t}_{n}$, $r^{t}_{n}$ is the return of the stock, and $\sum^{N}_{n=1}w^{t}_{n}=1$. To determine the weight of each stock, a trading strategy $\Phi\in\mathbb{R}^{\varrho}$ and the decision factor $\mathbf{b}$ is used, where $\mathbf{b}$ determines how many quantile groups are selected, satisfying $1\leq\mathbf{b}<\varrho$. The strategy does not allow short selling (as the Chinese stock market does not allow short selling), so $\Phi$ only contains 0 or $1$.
[…]
For example:
  • If $\varrho=3$, $\mathbf{b}=1$, then $\boldsymbol{\Phi}$ could be one of: $[1,0,0]$, $[0,1,0]$, or $[0,0,1]$.
  • If $\mathbf{b}=2$, then possible values of $\boldsymbol{\Phi}$ include: $[1,1,0]$, $[1,0,1]$, or $[0,1,1]$.
Recall the predicted output $\hat{y}^{t}_{n}=\{\hat{y}^{t}_{n,i}\}^{\varrho}_{i=1}$, we sort $S^{t}$ by $\Psi(\hat{y}^{t}_{n,1})$, which is the empirical quantile CDF for $\hat{y}^{t}_{n,1}$. Similar to the equation 6, $\Psi(\hat{y}^{t}_{n,1})$ is computed in this way:
$\Psi(\hat{y}^{t}_{n,1})=\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}_{\{\hat{y}^{t}_{i,1}\leq\hat{y}^{t}_{n,1}\}}$ (15)
The sorted predicted vector $\tilde{y}^{t}_{n}$ for stock $s^{t}_{n}$ at time t is shown below. Here the stocks are ranked into $\varrho$ parts and each part contains $\varphi\times 100\%$ of the stocks, and the parameters $\varrho$ and $\varphi$ satisfy $\varphi\varrho\leq 1$:
$\tilde{y}^{t}_{n}=[\mathbbm{1}_{\{(i-1)(\varrho+\xi)\leq\Psi(\hat{y}^{t}_{n,1})<i\varrho\ +(i-1)\xi)}|i=1,\dots,\varrho]\in\mathbb{R}^{\varrho},\quad\xi=\frac{1-\varphi\varrho}{\varrho-1}$
Then the weight for each selected stock is computed using equation 16 shown below, and all the chosen stocks are equal-weighted in this way:
$w^{t}_{n}=\frac{1}{\mathbf{b}\cdot\varrho}(\tilde{y}^{t-1}_{n}\Phi^{T}\tilde{y}^{t}_{n})^{T}\cdot\mathbf{1}$ (16)
The same method is run repeatedly during the subsequent periods. The backtest starts from January 2020, in other words, the result of the sequences from May 2018 to December 2019 will be used as the first stock pool to trade. The exchange fee in the Chinese stock market is 0.00384% (Shanghai stock exchange, 2023; Shenzhen stock exchange, 2023), and the brokerage commission is capped at 0.3% of the transaction amount (China securities regulatory commission, 2002). To keep our backtest results conservative, we set the transaction fee in the backtest to 0.3%.
@@
1:The feature sequence $\chi^{t}_{n}$ for each stock $s^{t}_{n}\in S^{t}$, initial portfolio cash $P^{0}$, trading strategy $\Phi$, strategy decision factor $\mathbf{b}$.
2:run_frequency(trade, frequency, startday=1, time=‘open’)
3:function Predict($\chi^{t}_{n}$): return $\hat{y}^{t}_{n}=\text{quantformer}(\chi^{t}_{n})$
4:end function
5:function SortLabel($\hat{y}^{t}_{n,1}$): return $\tilde{y}^{t}_{n}=\left[\mathbbm{1}\!\left\{\frac{i-1}{\varrho}\leq\Psi(\hat{y}^{t}_{n,1})<\frac{i}{\varrho}\right\}\right]_{i=1}^{\varrho}$
@@
7:function ComputeWeight($\tilde{y}^{t-1}_{n}$, $\tilde{y}^{t}_{n}$, $\Phi$):
8:  numerator = $\tilde{y}^{t-1}_{n}\cdot\Phi^{\top}\cdot\tilde{y}^{t}_{n}$
9:  denominator $\sum_{k=1}^{N}\tilde{y}^{t-1}_{k}\cdot\Phi^{\top}\cdot\tilde{y}^{t}_{k}$
10:  return $w^{t}_{n}=\frac{1}{\mathbf{b}\cdot\varrho}\cdot\frac{\text{numerator}}{\text{denominator}}$
11:end function
12:function Trade($w^{t}_{n}$, $P^{t}$):
13:  for $s^{t}_{n}$ in $S^{t}$ do order($s^{t}_{n}$, $w^{t}_{n}\cdot P^{t}$)
14:  end for
15:end function
@@ 4.5 Metrics
The Sharpe Ratio (SR) (Sharpe, 1964) and the $\alpha$ rate will be used to test the performance of the strategy. The SR is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase (equation (17)), where $R_{p}$ is the return of the portfolio and $R_{f}$ is the risk-free rate, chosen to be the London Interbank Offered Rate (LIBOR) which averages rate estimates submitted by banks in London [^3] Although LIBOR officially came to an end on September 30th, 2024, the time period covered by our backtesting dataset is prior to this date. https://www.bankofengland.co.uk/news/2024/october/the-end-of-libor. The risk-free rate in the backtest will be calculated as the average LIBOR rate during the test period.
@@
Alpha represents the excess return of an investment portfolio relative to its benchmark. It aims to measure the stock selection skills of an investment strategy. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment. Specifically, alpha is the excess return of the actual portfolio return over its expected theoretical return. Equation (18) shows the calculation method of the alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and the benchmark’s systemic risk.
$\alpha=(R_{p}-R_{f})-\beta(\mathbb{E}[R_{m}]-R_{f})$ (18)
Besides the SR ratio and alpha rate, the annual return (AR) of the stock, the annual excess return (AER), the win rate (WR) and the average turnover rate (TR) of the portfolio will also be shown. The excess returns are returns achieved above and beyond the return of a proxy, chosen to be the CSI 300 index. The win rate is the percentage of trading days during which the portfolio strategy generates a positive return. The turnover rate of the portfolio is the proportion of portfolio weight changes between two consecutive rebalancing periods. The definition of TR at each time $t$ is:
$\mathrm{TR}^{t}=\frac{1}{2}\sum^{N}_{k=1}|w^{t}_{n}-w^{t-1}_{n}|$ (19)
where $w^{t}_{n}$ and $w^{t-1}_{n}$ denote the portfolio weight of the stock $s_{n}$ in the portfolio at time $t$ and t1t-1, respectively.
Value at risk (VaR) is a method to summarize the total risk in a portfolio (Hull, 2012).
$\mathrm{VaR}_{\alpha}=\inf\{x:P(\mathcal{L}>x)\leq\alpha\}$ (20)
@@ 5 Results and discussion
Building upon the analysis of individual strategies, this section presents an evaluation and discussion of the experimental outcomes. The discussion focuses on the comparative effectiveness of different time frequencies (monthly, weekly, and daily) and different training scales.
@@ 5.1 Overall performance
Our results are shown in Table 3. For the monthly group, Month_1 shows an annual return of 17.35% and an annual excess return of 19.43%, indicating a robust performance over that period. Conversely, Month_3 shows a diminished annual return and annual excess return of 7.37% and 9.91% respectively. The is notably high in the second month at 51.69%. The win rate ranges from 43.1% to 57.3%, with the first month showing the highest win rate. The Sharpe Ratio and Alpha vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values. In the comparison of 99% VaR, the strategies in monthly frequency perform better than others, where Month_1 and Month_3 are lower than 3%. This means that, with 99% confidence, the portfolio maximum daily losses are 2.81% and 2.3% respectively.
@@
In the weekly strategy results, Week_3 shows an annual return of 12.3% and an annual excess return of 12.73%, while Week_1 displays a negative annual return. For the daily strategies, Day_2 has a return of 10.23% and an annual excess return of 10.94%. The turnover rate across the daily strategies shows lesser variation than the monthly or weekly, indicating a more consistent trading frequency. The win rate for the daily strategies remain fairly stable, hovering around the mid-40% mark. The Sharpe Ratio and Alpha for weekly and daily strategies show a mix of positive and negative values, reflecting the fluctuating nature of shorter-term trading efficacy.
To demonstrate the advantages of the quantformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^4] https://www.joinquant.com/ (one of China’s largest quantitative finance platforms) are selected for backtesting under the same trading strategy (Algorithm 1). Price-volume factors are used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by quantformer factors. The detailed performance of factors and benchmark is shown in A. On average, these factors achieves an annual return of -3.78%, an average excess return of -2.15%, and a Sharpe ratio of -0.36, where the benchmark return is -1.77%. The average maximum drawdown of the strategy is 44.88%, and the average volatility of the strategy ($\sigma_{p}$) is 0.26, with the benchmark at 0.197.
@@
Among these factor-based strategies, the quantformer factor performs well. Except for the win rate, which is slightly lower than a few strategies, it performs well in terms of annual return, annual excess return, $\sigma_{p}$, Sharpe ratio and Sortino ratio, ranking the best among the 101 factors. The 99% VaR of the quantformer factor is among the best 10% factors. Figure 4 illustrates the return curves of the quantformer (QF_Month_1) alongside the benchmark and some traditional factors. “25_week_rank” represents the current price’s position over the past 25 weeks; “EMAC20” and “MAC20” are the 20-day index moving average and the stock’s 20-day moving average respectively; “DAVOL20” and “VOL20” represent the ratio of the 20-day average turnover rate to the 120-day average turnover rate, and the mean of the stock’s 20-day turnover rate; “ROC20” is the price rate of change over 20 days; “MAWVAD” is calculated as the product of the difference between the closing price and the opening price, divided by the range of the highest and lowest prices, all multiplied by the volume, accumulated over six days; “Variance20” is the variance of the stock’s 20-day annualized returns. Most of these factors are compared based on a window of 20 timestamps, as the quantformer factor was also trained on 20 timestamps. In comparison with these factors, the quantformer factor (blue line) demonstrates a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies. Compared to other factors by the Sharpe ratio, the quantformer (0.915) still performs better than others (the highest Sharpe ratio among these factors is 0.243). Furthermore, the quantformer also shows a lower downside risk in terms of risk assessment as it achieves a VaR of 2.81, compared to the highest VaR of 2.86 among other factors. These indicate the improvement of the quantformer factor over traditional price-volume factors.